OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition?

OpenStudy (anonymous):

Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1 help me plzzzzzzz.

OpenStudy (anonymous):

please i need help this is my last try on this quiz

OpenStudy (anonymous):

i will medal and fan